OpenStudy (anonymous):

Find an equation for the contour of f(x,y)=2x^2y+7x+10 that goes through the point (1,3).

Miracrown (miracrown):

Just plug in the point into that equation and then you're going to set 2x^2y+7x+10 equal to that number for the equation of the contour. Go ahead and start by plugging in(1,3) into that equation

OpenStudy (anonymous):

2(1)^2(3) + 7(1) + 10 = 2(3) + 7 + 10 = 6 + 7 + 10 = 23 @Miracrown

Miracrown (miracrown):

Great! So you set the expression equal to that number 2x^2y + 7x + 10 = 23 And that's the equation of your contour :)

OpenStudy (anonymous):

hmm wow thats easy, I don't really understand it but, ty so much :)

OpenStudy (anonymous):

that was correct :)

Miracrown (miracrown):

Ahaha! I'm glad, it looked hard from the outside but the inside was simple as pie ;)